# The scalar-torsion gravity corrections in the first-order inflationary models

### v1

@@ Abstract
Abstract: In this work, we consider the corrections to the cosmological models based on the teleparallel equivalent of general relativity induced by the non-minimal coupling between scalar field and torsion. To determine these corrections, we consider a power-law parametrization of the deviations between the teleparallel equivalent of general relativity and the scalar-torsion gravity. The estimates of the possible influence of the non-minimal coupling of the scalar field and torsion on the background cosmological parameters and the parameters of cosmological perturbations for the inflationary models implying linear relation between tensor-to-scalar ratio and spectral index of the scalar perturbations are obtained. A procedure for verifying these inflationary models based on observational constraints on the values of cosmological perturbation parameters is also considered.
@@ I Introduction
In modern cosmological models, an important role is played by the description of two stages of the accelerated expansion of the universe, namely the primary inflation of the early universe [1, 2] and the second inflation in the modern era [3] of the evolution of the universe. To describe the cosmological dynamics of the universe at the different stages of its evolution and observed large-scale structure of the universe, general relativity [1, 2] and various modified theories of gravity are considered [4, 5, 6, 7]. One such approach is the use of the teleparallel equivalent of general relativity (TEGR) to describe gravity, where torsion scalar $T$ is considered instead of the scalar curvature $R$ [8, 9]. Also, the different modifications of TEGR are considered to construct and analyze the cosmological models  [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
@@
Equations of cosmological dynamics for the case of $f(T)$-gravity is the second order as in the case of general relativity, however, instead of using the curvature defined through the Levi-Civita connection, in the case of $f(T)$-gravity the Weitzenböck connection is used, which implies zero curvature and non-zero torsion [25, 26, 27]. The fact that the dynamic equations for $f(T)$-gravity are second order [25] makes these theories simpler than those modified by curvature invariants (see, for example, in [4, 5, 6, 7]), and a study of this type of model is of considerable interest for theoretical cosmology. It should also be noted that the speed of gravitational waves for the case of $f(T)$-gravity is equal to the speed of light in vacuum, which is similar to both the case of general relativity and scalar-tensor theories of gravity and satisfies the restrictions on the speed of gravitational waves from astrophysical objects according to modern observations [28].
@@
Thus, the teleparallel equivalent of general relativity (TEGR) and its modifications are the basis for the construction of cosmological models which can be verifiable by observational data, that corresponds to the growing interest in this type of gravity models and their physical applications in modern cosmology  [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. When constructing and analyzing cosmological models based on TEGR modifications, the following questions arises: how do these modifications affect cosmological parameters and what possible physical effects do they correspond to? To answer these questions, the approach associated with a model-independent assessment of the influence of modified theories of gravity on cosmological parameters seems promising. In  [29, 30, 31, 32, 33, 34, 35] methods were proposed for analyzing cosmological models with modified gravity theories (scalar-tensor gravity and Einstein-Gauss-Bonnet gravity) based on functional and parametric connections to the Einstein gravity in the Friedman-Robertson-Walker space-time. This approach made it possible to directly assess the influence of these modifications of the Einstein gravity on the background dynamics and parameters of cosmological perturbations. In  [36] this approach was used for the reconstruction of the scalar-torsion gravity theories from the physical potentials of a scalar field.
@@
In this work, we present the model-independent approach to estimate the influence of non-minimal coupling between a scalar field and torsion on the cosmological parameters. In the context of the proposed approach to this interpretation of the description of the inflationary scenario, we consider the influence of the non-minimal coupling of the scalar field and torsion on the cosmological dynamics, the scalar field potential and the parameters of cosmological perturbations. The proposed power-law relationship between the coupling function and the Hubble parameter allows us to reduce the assessment of such influence to the estimation of the possible values of one constant parameter. It was also shown, that the corrections inspired by non-minimal coupling of the scalar field and torsion lead to the verification of the inflationary models implying a linear relation between tensor-to-scalar ratio and spectral index of the scalar perturbations by the observational constraints in contrast to the case of TEGR.
@@
The paper is organized as follows. In Section II we consider dynamical equations for cosmological models based on the scalar-torsion gravity. Section III proposes a power-law relation between the non-minimal coupling function and the Hubble parameter $F=(H/\lambda)^{n}$. The implications of this relation are also discussed. In Section IV, we obtain the parameter $\lambda$ in explicit form based on the conditions for the correct completion of the inflationary stage. Section V provides restrictions on the values of parameter $n$ and on the possible types of inflationary models. In Section VI it was shown that the considered inflationary models with non-minimal coupling between scalar field and torsion can be verified from observational data, unlike models with minimal coupling. Estimates of the propagation velocity of scalar perturbations for the verified inflationary models were also given. A summary of our investigations is presented in Sec. VII.
@@ II Dynamical equations for cosmological models based on the scalar-torsion gravity
The action of generalized scalar-torsion theory with self-interaction of Galilean-like field in the system of units $M^{2}_{p}=(8\pi G)^{-1}=c=1$ can be defined as follows [18]
@@
where $F(\phi)$ is the function of the scalar field $\phi$, $P$, $G$ are functions of the scalar field and its kinetic energy $X=-\frac{1}{2}\phi_{,\mu}\phi^{,\mu}$ and $\phi_{,\mu}=\partial_{\mu}\phi=\frac{\partial\phi}{\partial x^{\mu}}$, $T$ is the torsion scalar, $e^{A}_{\mu}$ are the tetrad components in a coordinate basis, $e=det(e^{A}_{\mu})=\sqrt{-g}$ and $G(\phi,X)$ is the Galileon-type field self-interaction. Also, the function $P$ can be considered as the function of the potential of a scalar field and its kinetic energy $P=P\left(V(\phi),X\right)$.
In this paper, we consider a special case of negligible self-interaction of a scalar field
@@
In this case, the action Eq. (1) is reduced to the following form
@@
Thus, we will consider the nonlinear self-interaction of the scalar field as a higher-order effect compared to the non-minimal coupling of the scalar field and torsion.
The tetrad field
@@
corresponds to the homogeneous and isotropic Friedman-Robertson-Walker (FRW) universe with metric
@@
where $a=a(t)$ is the scale factor which dependent on cosmic time $t$ for expanding universe.
Thus, the background equations for the action Eq. (3) and tetrad field (4) are [18, 19, 20]
@@
where overdot means the derivative with respect to cosmic time.
Let us consider a particular case where the function $P=P(\phi,X)$ is defined as follows
@@
where $\omega$ is a some constant.
Thus, from Eqs. (7)–(9) we obtain
@@
$\omega\dot{\phi}^{2}=-2H\dot{F}-2F\dot{H}.$ (11)
For the case of minimal coupling $F=1$, these models are reduced to the case of the inflationary models based on the teleparallel equivalent of general relativity.
@@ III Power-law parametrization of coupling function
In order to estimate the influence of the non-minimal coupling on the cosmological parameters, we consider some connection between coupling function $F=F(\phi)$ and the Hubble parameter $H=H(t)$ to parameterize the influence of the non-minimal coupling between the scalar field and torsion on the parameters of cosmological models under consideration.
We define this connection as the power-law one, namely
@@
where $\lambda$ is a positive constant and $n$ is the parameter that defines the influence of the non-minimal coupling between the scalar field and torsion. For $n=0$ one has the case $F=1$ corresponding to minimal coupling or TEGR.
@@
Dynamic equations (10)–(11), in this case, have the following form
@@
$\omega\dot{\phi}^{2}=-2\lambda^{-n}\dot{H}H^{n}(n+1).$ (14)
In this work, unlike paper [36], where kinetic function was considered as $\omega=\omega(\phi)$, we will consider $\omega=const$ corresponding to the canonical (for $\omega>0$) and phantom (for $\omega<0$) scalar fields.
Thus, equations (13)–(14) can be rewritten as
@@
where we will consider
$\omega=\lambda^{-n}=const,$ (17)
as a convenient normalization of the constant $\omega$ implying canonical scalar field $\omega=1$ for the minimal coupling $n=0$.
Thus, for the case of the minimal coupling $n=0$, dynamic equations (III)–(16) are reduced to the well-known expressions [2]
@@
The quasi de-Sitter expansion of the early universe $h\approx const$ implies the following slow-roll condition $X=\frac{1}{2}\dot{\phi}^{2}\ll V$ at the inflationary stage.
Let us introduce the connection between $H$ and $h$ like follow
@@
we can rewrite equations (III)–(19) as follows
$V(\phi)=\lambda^{-n}\left[3h^{\frac{2+n}{1+n}}+\dot{h}\right]=\lambda^{-n}\left[3h^{\frac{2+n}{1+n}}+\frac{1}{2}\dot{\phi}^{2}\right],$ (22)
@@
Also, using Ivanov-Salopek-Bond approach [2] based on the relation $\dot{h}=\dot{\phi}h_{,\phi}=\dot{\phi}\frac{dh(\phi)}{d\phi}$, we can rewrite (22)–(23) as
@@
Also, we note, that under condition $01+n>$ 0 for $\dot{h}<0$ one has $\dot{H}<0$ as well. Thus, under condition $01+n>$ 0 the inflationary stage is being implemented for the case of non-minimal coupling if inflation occurs for the case of minimal coupling.
The condition of inflationary stage $01+n>$ 0 restricts the possible values of the characterization constant as follows
@@
Also, under the slow-roll conditions $\tilde{X}\ll\tilde{V}$ and $X\ll V$, from expressions (III), (18) and (24) we get
@@
Thus, from (28)–(29) we obtain the following relation between potentials for non-minimal and minimal coupling
@@
Also, from expressions (27) and (28) we obtain the relation between the coupling function and the potential for minimal coupling
@@
Further, based on relations (20)–(21), we will express the slow-roll parameters for the case of non-minimal coupling in terms of the Hubble parameter $h=h(t)$ as follows
@@
For the case $n=0$, from (32)–(33) one has the slow-roll parameters for inflation with minimal coupling [1, 2]
@@
IV State parameter at the end of the inflationary stage
Now, we determine the constant parameter $\lambda$ from the conditions on the state parameter at the end of the stage of cosmological inflation. In cosmological models with minimal coupling, the state parameter of a scalar field is determinate as follows [1, 2]

### v3

@@ Abstract
Abstract: The corrections to the cosmological models induced by non-minimal coupling between scalar field and torsion are considered. To determine these corrections in explicit form, the power-law parametrization of these corrections are proposed. The estimates of possible influence of non-minimal coupling between scalar field and torsion on cosmological parameters for inflationary models implying linear relation between tensor-to-scalar ratio and spectral index of scalar perturbations are obtained. A procedure for verifying these inflationary models due to the observational constraints on the values of cosmological perturbation parameters is also considered.
@@ I Introduction
In the current cosmological models, a crucial aspect is the characterization of two phases of accelerated expansion: the initial inflationary era in the early Universe [1, 2] and the accelerated period in the present cosmic era [3]. To describe the dynamics of the expansion of the Universe at different stages of its evolution and observed large-scale structure of the Universe, general relativity [1, 2] and various modified theories of gravity are considered [4, 5, 6, 7]. Teleparallel equivalent of general relativity (TEGR) is another geometrical approach, where torsion scalar $T$ is considered instead of scalar curvature $R$ [8, 9, 10, 11]. The different modifications of TEGR are considered to construct and analyze the cosmological models both at the level of background dynamics [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29] and with taking into account evolution of cosmological perturbations [30, 31, 32, 33, 34, 35, 36] as well. We can also note the possibility of constructing the cosmological models based on braneworld teleparallel gravity  [37, 37, 38, 39, 40, 41]. It should be noted that cosmological models including scalar field as the source of accelerated expansion of the Universe can be considered with taking into account the Galilean field self-interaction [33] and without this term [34] as well.
@@
In $f(T)$ gravity, equations of cosmological dynamics are of second order as in the case of general relativity(GR). However, instead of using the curvature defined through the Levi-Civita connection, in case of $f(T)$ gravity the Weitzenböck connection is used, which implies zero curvature and non-zero torsion [42, 43, 44]. The fact that the dynamic equations for $f(T)$ gravity are of second order [42], it makes these theories simpler than those modified by curvature invariants (see, for example, in [4, 5, 6, 7]). This type of cosmological model is of considerable interest in theoretical cosmology. It should also be noted that the speed of gravitational waves for the case of $f(T)$ gravity is equal to the speed of light in vacuum, which is similar to both the case of GR and scalar-tensor theories of gravity and satisfies the restrictions on the speed of gravitational waves from astrophysical objects according to the current cosmological observations [45].
It should also be noted that scalar-torsion gravity theories are also used to construct cosmological models of the current stage of the evolution of the Universe [24, 25, 26, 27, 28, 46]. Currently, various possibilities for modifying the standard $\Lambda$CDM-model are being considered to alleviate $H_{0}$ and $\sigma_{8}$ tension in the $\Lambda$CDM-model (see, Refs. [47, 48, 49, 50, 51, 52, 53, 54, 55]). The models based on the scalar-torsion gravity including the Galilean field self-interaction [28] and without this term [46] can be considered to modify the $\Lambda$CDM-model.
@@
Thus, the teleparallel equivalent of general relativity (TEGR) and its modifications are the basis for the construction of cosmological models which can be verifiable by observational data, that corresponds to the growing interest in this type of gravity models and their physical applications in current cosmology  [15, 16, 17, 18, 19, 30, 31, 32, 33, 34, 20, 21, 22, 23, 24]. When constructing and analyzing cosmological models based on TEGR modifications, the following questions arise: how do these modifications affect cosmological parameters and what possible physical effects do they correspond to? To answer these questions, the approach associated with a model-independent assessment of the influence of modified theories of gravity on cosmological parameters seems promising. In [56, 57, 58, 59, 60, 61, 62, 63] methods were proposed for analyzing cosmological models with modified gravity theories (scalar-tensor gravity and Einstein-Gauss-Bonnet gravity) based on functional and parametric connections to the Einstein gravity in the Friedman-Robertson-Walker space-time. This approach made it possible to directly assess the influence of these modifications of the Einstein gravity on the background dynamics and parameters of cosmological perturbations. In [35] this approach was used for the reconstruction of the scalar-torsion gravity theories from the physical potentials of a scalar field. Also, in [29, 36] the cosmological models
@@
In this work, we present a model-independent approach to assess the impact of the non-minimal coupling between the scalar field and torsion on cosmological parameters during the inflationary phase of the Universe evolution. In this approach, the influence can be estimated explicitly for cosmological dynamics, scalar field potential and cosmological perturbations parameters. The proposed power-law relationship between coupling function and Hubble parameter allows us to reduce the assessment of such influence to the estimation of the possible values of the velocity of the scalar perturbation. This feature of the proposed approach distinguishes it from previously considered methods of analyzing cosmological inflation models based on scalar-torsion gravity theories (see, Refs. [33, 34, 20, 35, 29, 36]). The corrections inspired by non-minimal coupling between scalar field and torsion lead to the verification of the inflationary models implying a linear relation between tensor-to-scalar ratio and spectral tilt of scalar perturbations by the observational constraints.
@@
The paper is organized as follows: In Sec-II, we have considered the dynamical equations for cosmological models based on the scalar-torsion gravity. Sec-III proposes a power-law relation between the non-minimal coupling function and the Hubble parameter $F=(H/\lambda)^{n}$. The implications of this relation are also discussed. In Sec-IV, we have obtained the parameter $\lambda$ in explicit form based on the conditions for the correct completion of the inflationary stage. Sec-V provides restrictions on the values of parameter $n$ and on the possible types of inflationary models. Also, it was shown that reducing the cosmological models under consideration to the case of $f(T)$-gravity leads to the pure de Sitter stage. In Sec-VI we formulate conditions on the slow-roll parameters and condition of the exit from inflation, which restrict the possible types of inflationary models. In Sec-VII, it has been shown that the considered inflationary models with non-minimal coupling between scalar field and torsion can be verified from observational data, unlike models with minimal coupling. Estimates of the propagation velocity of scalar perturbations for the verified inflationary models are also given. A summary of our investigations has been presented in Sec. VIII.
@@ II Dynamical equations for cosmological models based on the scalar-torsion gravity
The action of generalized scalar-torsion theory with self-interaction of Galilean-like field in the system of units $M^{2}_{p}=(8\pi G)^{-1}=c=1$ can be defined as [33],
@@
where $F(\phi)$ is the function of the scalar field $\phi$; $P(\phi,X)$, $G(\phi,X)$ are functions of the scalar field and its kinetic energy $X=-\frac{1}{2}\phi_{,\mu}\phi^{,\mu}$ and $\phi_{,\mu}=\partial_{\mu}\phi=\frac{\partial\phi}{\partial x^{\mu}}$. $T$ be the torsion scalar, $e^{A}_{\mu}$ are the tetrad components in a coordinate basis, $e=det(e^{A}_{\mu})=\sqrt{-g}$ and $G(\phi,X)$ is the Galileon-type field self-interaction. Also, the function $P$ can be considered as the function of the potential of a scalar field and its kinetic energy $P=P\left(V(\phi),X\right)$.
The torsion scalar is defined as follows [33, 34, 20]
$T=S_{\gamma}^{~{}\alpha\beta}\,T^{\gamma}_{~{}\alpha\beta}\,,$ (2)
$T^{\gamma}_{~{}\alpha\beta}=e_{A}^{~{}\gamma}\left[\partial_{\alpha}e^{A}_{~{}\beta}-\partial_{~{}\alpha}e^{A}_{~{}\beta}+\omega^{A}_{~{}B\alpha}\,e^{B}_{~{}\beta}-\omega^{A}_{~{}B\beta}\,e^{B}_{~{}\alpha}\right]\,,$ (3)
$S_{\gamma}^{~{}\alpha\beta}=\frac{1}{2}\left(K^{\alpha\beta}_{~{}~{}\gamma}+\delta^{\alpha}_{~{}\gamma}\,T^{\mu\beta}_{~{}~{}\mu}-\delta^{\beta}_{~{}\gamma}\,T^{\mu\alpha}_{~{}~{}\mu}\right)\,,$ (4)
$K^{\alpha\beta}_{~{}~{}\gamma}=-\frac{1}{2}\left(T^{\alpha\beta}_{~{}~{}\gamma}-T^{\alpha\beta}_{~{}~{}\gamma}-T_{\gamma}^{~{}\alpha\beta}\right)\,,$ (5)
where $T^{\gamma}_{~{}\alpha\beta}$ is the torsion tensor, $S_{\gamma}^{~{}\alpha\beta}$ is the super-potential, $K^{\alpha\beta}_{~{}~{}\gamma}$ is the contorsion tensor, and spin connection $\omega^{A}_{~{}B\alpha}=0$ for a special class of inertial frames.
We consider a special case of negligible self-interaction of a scalar field as,
@@
So, the action Eq. (1) is reduced to the following form
@@
In [33], the different inflationary models corresponding to assumption (6) and with non-negligible Galileon field self-interaction were considered. Also, the verifiable models of the current stage of the Universe’s evolution can be constructed with non-negligible Galileon self-interaction [28] and without this term [46] as well. Thus, we will use assumption (6) to consider the need to include the Galilean field self-interaction in actual cosmological models.
Thus, we shall consider the nonlinear self-interaction of the scalar field as a higher-order effect compared to the non-minimal coupling of the scalar field and torsion. The tetrad field
@@
corresponds to isotropic and homogeneous Friedmann-Lemaitre-Robertson-Walker (FLRW) metric
@@
where $a=a(t)$ is the scale factor that dependents on cosmic time $t$ only.
For FLRW metric (8)–(9) expressions (2)–(5) imply following expression for torsion scalar
$T=6H^{2}.$ (10)
Background equations for action Eq. (7) and tetrad field (8) are [33, 34, 20]
@@
where an over dot denotes ordinary derivative concerning cosmic time $t$.
For brevity, we denote $P\approx P(\phi,X)$ and consider
@@
where $\omega$ is a constant.
From Eqs. (12)–(14), we obtain
@@
$\omega\dot{\phi}^{2}=-2H\dot{F}-2F\dot{H},$ (16)
$\omega\ddot{\phi}+3\omega H\dot{\phi}+V_{,\phi}+3H^{2}F_{,\phi}=0.$ (17)
To note, for the case of minimal coupling ($F=1$), these models are reduced to the case of inflationary models based on TEGR.
We also note that of the three equations of cosmological dynamics (15)–(17) only two are independent. Thus, we can use any two equations in system (15)–(17) to analyze the models of cosmological inflation.
@@ III Power-law parametrization of coupling function
To estimate the influence of non-minimal coupling on the cosmological parameters, we consider some connection between coupling function $F=F(\phi)$ and Hubble parameter $H=H(t)$. We define the power-law connection as,
@@
where $\lambda$ is a positive constant and $n$ is the parameter that defines the influence of the non-minimal coupling between the scalar field and torsion.
For $n=0$ one has $F=1$, and action (7) is reduced to following form
$S\approx\int d^{4}xe\left[\frac{1}{2}T+P(\phi,X)\right],$ (19)
corresponding to the case of the teleparallel equivalent of general relativity [33, 34, 20].
The power-law connection between coupling function and Hubble parameter was considered earlier in [56, 57, 58] in the context of the model-independent approach to verification of the inflationary models based on the scalar-tensor gravity theories. Also, relation (18) was considered in [35] for the case of the inflationary models based on the scalar-torsion gravity with arbitrary kinetic function $\omega=\omega(\phi)$. In [35] the constant $n$ characterized the influence of the non-minimal coupling of the scalar field and torsion on cosmological parameters, but it was not itself directly related to any cosmological parameter. In the framework of the new approach, we will relate both free constant parameters in relationship (18) with the speed of propagation of scalar perturbations. This method of parameterizing the influence of the scalar-torsion modifications of TEGR leads to additional physical motivated constraints on the constants $n$ and $\lambda$ in relationship (18) compared to the approach proposed in [35].
@@
Now, Eqs.(15)–(16) become
@@
$\omega\dot{\phi}^{2}=-2\lambda^{-n}\dot{H}H^{n}(n+1),$ (21)
where $\omega=constant$ corresponding to the canonical (for $\omega>0$) and phantom (for $\omega<0$) scalar fields.
Thus, Eqs. (20)–(21) can be rewritten as,
@@
where
$\omega=\lambda^{-n}=constant.$ (24)
The convenient normalization of the constant $\omega$ implying canonical scalar field $\omega=1$ for the minimal coupling $n=0$. To note in minimal coupling, dynamic equations (III)–(23) are reduced to the well-known expressions [2]
@@
The quasi de-Sitter expansion of the early Universe $h\approx constant$ implies the following slow-roll condition $X=\frac{1}{2}\dot{\phi}^{2}\ll V$ at the inflationary stage.
We introduce the connection between $H$ and $h$ as,
@@
Eqs. (III)–(26) can now be rewritten as,
$V(\phi)=\lambda^{-n}\left[3h^{\frac{2+n}{1+n}}+\dot{h}\right]=\lambda^{-n}\left[3h^{\frac{2+n}{1+n}}-\frac{1}{2}\dot{\phi}^{2}\right],$ (29)
@@
Further, using Ivanov-Salopek-Bond approach [2] based on the relation $\dot{h}=\dot{\phi}h_{,\phi}=\dot{\phi}\frac{dh(\phi)}{d\phi}$, one can rewrite Eqs. (29)–(30) as
@@
Under condition $01+n>$ 0 for $\dot{h}<0$ one can have $\dot{H}<0$ as well. So, under condition $01+n>$ 0 the inflationary stage is being implemented for the case of non-minimal coupling if inflation occurs for the case of minimal coupling. The condition of inflationary stage $01+n>$ 0 restricts the possible values of the characterization constant as,
@@
Under the slow-roll conditions $\tilde{X}\ll\tilde{V}$ and $X\ll V$, from expressions (III), (25) and (31) we get
@@
Thus, from Eqs. (35)–(36) we obtain the following relation between potentials for non-minimal and minimal coupling
@@
Also, from Eqs. (34) and (35), we can obtain the relation between the coupling function and the potential for minimal coupling
@@
Further, based on relations (27)–(28), we will express the slow-roll parameters for the case of non-minimal coupling in terms of the Hubble parameter $h=h(t)$ as,
@@
For the case $n=0$, from Eqs. (39)–(40) one can have the slow-roll parameters for inflation with minimal coupling [1, 2]
@@
IV Equation of state (EoS) parameter at the end of the inflationary stage
Now, we shall determine the constant parameter $\lambda$ from the conditions on the EoS parameter at the end of the stage of cosmological inflation. In cosmological models with minimal coupling, the EoS parameter of a scalar field can be given as [1, 2],